# Learning Cross-Task Generalities Across Graphs via Task-trees

### v1

@@ 5.4 Comparison Between Task-Trees and Subgraphs
To illustrate this, we compare the efficiency and performance of subgraphs and task-trees in the basic fine-tuning setting, as shown in Figure 6 and Table 5, respectively. We implement a subgraph version of GIT by replacing task-trees with subgraphs (GIT-Subgraph). The results show that task-trees offer both better efficiency and improved model performance. Although Table 5 presents results from only three domains, the evaluation covers a range of tasks, including node classification (academic networks), edge classification (knowledge graphs), and graph classification (molecule networks).
6 Conclusion
Conclusion. We introduce the concept of task-trees as basic learning instances for graphs and provide both theoretical and empirical validation of the task-tree generality assumption: the generalities shared across graphs are (at least partially) preserved in the task-trees of the involved graphs. Building on task-trees, we develop GIT, a cross-domain and cross-task GFM that can be quickly specialized for specific domains. By pretraining on a small set of graphs, GIT improves performance across more than 30 graphs spanning five domains. After specialization with a small set of domain-specific task-trees, GIT can match or even surpass expert models designed for specific domains.
Limitations and Future Work. We treat task-trees as basic learning instances in graphs, similar to how images are used in vision and text in language. Despite we have validated that cross-graph generalities are, at least partially, preserved in task-trees, identifying the exact transferable patterns shared across graphs remains a challenge. Further exploration of this issue from both empirical and theoretical perspectives is worthwhile. Additionally, for message-passing GNNs, converting graphs into tree structures inevitably leads to some loss of information, limiting the model expressiveness. Advanced GNNs (Morris et al., 2019; 2023) or techniques (Zhang et al., 2021) could be used to enhance expressiveness. Moreover, as GIT is a relatively simple model, investigating potential extensions is also a promising direction, as discussed in Appendix B.
@@ 6 Conclusion
Broader Impact. From an industry perspective, we offer GIT as a foundational tool for graph-structured data. Additionally, since GIT can be quickly adapted to specific domains, we hope it will support applications where label acquisition is difficult and model training is time-consuming.
Appendix: Table of Content
[…]
@@ Appendix A Related Work
Graph Foundation Models. Foundation models are designed as general-purpose solvers capable of handling various tasks across different domains. For instance, LLMs, the foundation models in natural language processing, are capable of performing tasks such as summarization, translation, and entity recognition, as well as question-answering. However, building such versatile foundation models for graphs presents unique challenges due to the inherent feature, structural, and task heterogeneity across different graph domains and tasks. To address these challenges, Qiu et al. (2020) pretrained GNNs using subgraphs as basic units, mitigating structural heterogeneity. Building on this, Sun et al. (2023) reformulated node-, edge-, and graph-level tasks into subgraph-level tasks, tackling task heterogeneity. Additionally, Huang et al. (2023) and Liu et al. (2024) applied LLMs to unify the feature spaces of cross-domain graphs, addressing feature heterogeneity. These approaches enable models to operate on cross-domain and cross-task graphs. Further advancements, such as He & Hooi (2024) and Li et al. (2024), improve node embeddings by jointly optimizing GNN and LLM encoders, facilitating various downstream tasks like few-shot learning and zero-shot learning. Other efforts to resolve feature heterogeneity include methods like singular vector decomposition (SVD) (Zhao et al., 2024a; Yu et al., 2024) and non-parametric encoders (Zhao et al., 2024b). However, most of these approaches rely on subgraphs as the primary learning instances, which can result in inefficient training and reduced expressiveness, as discussed in the main paper.
@@
Appendix B Potential Model Extensions
B.1 Pretraining
B.2 Specialization via Instruction Tuning
How to Define Instructions? In this paper, as we focus on experiments with text-attributed graphs Wang et al. (2024b), we define instructions as label descriptions encoded by LLMs. However, this approach is not applicable to non-textual graphs. Other methods could be explored to define instructions, such as using proxy models (Hu et al., 2019) or graph heuristics (Jin et al., 2020) to generate instructions.
How to Leverage SFT Data? In scenarios with limited instructions, standard supervised fine-tuning may struggle to capture sufficient knowledge of the target domain. To address this, methods could be proposed to better utilize the unlabeled instances in the SFT dataset, thus enhancing model adaptation (Sohn et al., 2020; Wang et al., 2025).
B.3 More Scenarios.
The paper leverages text-attributed graphs to align node features. However, the pre-processing of TAGs can be time-consuming, raising the challenge of how to effectively apply the model to graphs without aligned node features. Furthermore, while we primarily focus on homogeneous graphs in this work, most real-world applications involve heterogeneous graphs. Addressing the question of how to design a single model capable of handling various types of graphs remains an open challenge. Finally, applying the model to specific applications Zhang et al. (2024c; 2025), which may exhibit unique characteristics, is another important consideration for future research.
Appendix C Additional Discussion
C.1 Why Does the General Model Need Specialization?
@@
C.2 More Analysis on Domain Regularizer
((a))
((b))
((c))
((d))
@@ C.2 More Analysis on Domain Regularizer
C.3 Discussion on Homophily and Heterophily
@@
C.4 Discussion on Model Expressiveness
@@
C.5 Scaling Law
((a))
((b))
@@ C.5 Scaling Law
Model Size. We evaluated the model’s performance with different hidden dimensions, with results by domain presented in Figure 9. The results cover both basic fine-tuning and in-context learning, and comprehensive details are provided in Appendix F.8. We observe that increasing the number of hidden dimensions from 128 to 2,048 significantly improves model performance across all domains. We hypothesize that this improvement is due to the additional parameters, which enhance the model’s ability to memorize shared patterns across graphs. The observation indicates the potential existence of scaling laws when using task-trees as the basic learning instances.
@@ Appendix D Proof
By using a formulation like expression 10, we can decompose the impact of different layers, facilitating further analysis of the upper bound on the distance between two task-trees. Next, we will analyze the upper bound of each term. To begin, we first introduce a lemma.
@@ E.2 Baselines
Few-shot Learning Methods
GPN (Ding et al., 2020). GPN is a framework that leverages GNN and meta-learning to address few-shot node classification by learning a transferable metric space.
TENT (Wang et al., 2022b). TENT introduces three levels of adaptation—node-level, class-level, and task-level—to mitigate task variance and improve the model’s generalization performance across different meta-tasks.
GLITTER (Wang et al., 2022a). Enhance few-shot node classification by learning task-specific graph structures for each meta-task using node influence and mutual information.
TLP (Tan et al., 2022). transfer pretrained node embeddings fine-tunes a simple linear classifier on novel classes.
Prodigy (Huang et al., 2023). Enable in-context learning over graphs by designing graph prompt learning template.
@@ E.5 Hyper-Parameters
GIT. The model architecture and pretraining parameters of our GIT are presented in Table 8. The specific fine-tuning hyperparameters, categorized by domain, are shown in Tables 9, 10, 11, 12, and 13. For in-context learning and zero-shot learning results without fine-tuning, the general model does not involve any hyperparameters. For the specialized model, we tune the hyperparameters of SFT epochs from 10 to 500, in steps of 10, the SFT learning rate from 1e-4, 1e-5, 1e-6, 1e-7, 1e-8, and the normalization method from None, BN.
Appendix F Additional Results
F.1 Main Results
Figure 11: The model performance on all datasets in the zero-shot setting.
In-context Learning. The model performance in in-context learning, along with the best baselines, is shown in Figure 3. The results are presented from best to worst. It is evident that the specialized version significantly improves the general model, particularly in molecule and knowledge graphs. An interesting observation is that, while specialization enhances performance across datasets within the domain, it does not necessarily lead to a significant improvement on the dataset used for specialization. We hypothesize this is due to the use of prototypes for predictions, where the prototypes are constructed via random sampling, which may not be directly influenced by the supervised fine-tuning process. Detailed experimental results are presented in the following sections.
Zero-shot Learning. The zero-shot learning results across all datasets are shown in Figure 11. Compared to basic fine-tuning and in-context learning, the improvements in zero-shot performance are the most pronounced. Unlike the observations in in-context learning, the model performance on SFT datasets (Arxiv, Products, PCBA) is significantly higher. This is likely because the model is directly fine-tuned using label description embeddings, which are also used as prototypes during predictions. To reduce the influence of SFT data on zero-shot learning, we report the average performance on held-out graphs (i.e., graphs not used in pretraining or specialization) in Table 1 of the main paper.
F.2 Domain: Academia
@@
5-way 3-way
5-shot 3-shot 1-shot 5-shot 3-shot 1-shot
Comparing to Few-shot Learning Methods. We compare the performance of GIT with methods designed for few-shot learning on Arxiv in Table 16. It is important to note that we implement an in-context version of few-shot learning without fine-tuning GIT on the few-shot task, whereas the other methods require fine-tuning. Despite this, GIT achieves the second-best performance across all settings, outperforming 5 out of 6 few-shot learning methods. This highlights GIT’s strong potential in scenarios with limited instances and labels.
Ablation Study on Training Strategy. We perform an ablation study on academic networks to assess the impact of different training strategies. Specifically, we evaluate four strategies: (1) base model: pretraining and fine-tuning on the same graph, (2) domain expert: pretraining on all academic networks, (3) general model: pretraining on the default datasets used in this paper, and (4) specialized model: pretraining on the default datasets followed by specialization on the academic domain. The results, averaged across all academic graphs, are presented in Table 17. We make the following observations: (1) When pretrained on cross-domain and cross-task datasets (general model), GIT achieves the best performance, highlighting the advantages of using task-trees as the basic learning instance. However, the model does not consistently outperform in the base model and domain expert settings, possibly due to its model simplicity. (2) The general version of GIT maintains consistent performance compared to the base and expert models, while GraphMAE and OFA show a performance drop when transitioning from the base and expert models to the general model. This demonstrates the potential of task-trees in mitigating negative transfer in cross-domain and cross-task settings. (3) Interestingly, specialization enhances the performance of all baselines, showcasing the potential of instruction tuning to improve model capabilities in specific scenarios.
Table 17: Ablation study on the academic domain, evaluating different training strategies. We report the average performance across all academic graphs. The base model is pretrained on the target graph, the domain expert is pretrained on all academic graphs, the general model is pretrained on the default pretraining datasets, and the specialized model is fine-tuned via instruction tuning on the Arxiv dataset, based on the general model.
Base Model Domain Expert
0-shot 3-shot Finetune 0-shot 3-shot Finetune
GraphMAE 15.30 51.51 75.57 17.89 55.88 75.26
OFA 14.19 50.15 75.12 17.26 54.88 75.97
GIT 15.36 53.31 75.53 18.38 55.10 75.47
General Model Specialized Model
0-shot 3-shot Finetune 0-shot 3-shot Finetune
GraphMAE 15.42 49.25 73.81 20.31 51.21 74.05
OFA 13.98 45.93 72.18 20.05 46.87 73.04
GIT 14.88 54.00 75.82 23.45 55.18 75.88
((a))
((b))
((c))
F.3 Domain: E-Commerce
@@
F.4 Domain: Knowledge Base
@@
3-shot Finetune
Ultra Expert GIT Specialized GIT Ultra Expert GIT Specialized GIT
Average 68.53 67.72 67.80 76.26 76.03 76.15
Comparison to Domain Experts. In addition to comparing GIT to standard baselines applicable for all graphs, we also evaluate it against Ultra, a foundation model specifically designed for knowledge graphs. As a domain expert, Ultra is compared to Expert GIT (pretrained on all KGs) and Specialized GIT (pretrained on default datasets and fine-tuned on FB15K237), with the results presented in Table 21. We find that the two domain experts, Ultra and Expert GIT, achieve comparable performance, though Ultra significantly outperforms Expert GIT in certain settings. This may be due to Ultra learning more fine-grained relational information within KGs. Notably, Specialized GIT also performs comparably to both domain experts, highlighting the potential of specialization. We believe this is because the distributions of KGs are more similar to each other compared to graphs from other domains.
F.5 Domain: Molecule
@@
HIV BBBP BACE TOXCAST CYP450 TOX21 MUV PCBA Avg.
KVPLM 61.20 60.20 51.26 50.96 59.22 49.17 61.72 48.11 55.23
MoMu 50.26 49.81 66.56 52.38 57.98 57.57 60.51 51.50 55.82
Galactica-1.3B 33.85 53.94 56.48 51.23 46.86 49.46 57.15 52.02 50.12
GIMLET 66.24 59.39 69.57 59.04 71.25 61.19 64.39 62.11 64.15
GIT - G 56.76 54.76 33.66 51.55 63.21 56.83 53.71 56.25 53.34
GIT - S 66.14 62.16 52.27 58.30 69.75 63.45 65.32 65.26 62.83
indicates the results from paper (Zhao et al., 2023).
Table 24: The impact of SFT datasets in zero-shot setting.
SFT Data HIV BBBP BACE TOXCAST CYP450 TOX21 MUV PCBA Avg.
PCBA 66.14 62.16 52.27 58.30 69.75 63.45 65.32 65.26 62.83
HIV 66.28 45.97 43.35 52.78 64.50 57.86 53.46 46.57 53.85
Experimental Settings. We evaluate fine-tuning, in-context learning, and zero-shot learning in this domain. The fine-tuning and in-context learning settings are consistent with those used in previous domains. For zero-shot learning, however, we follow the approach of Zhao et al. (2023) by assessing zero-shot performance on the original test set.
Graph Classification. The graph classification results are presented in Table 22. Our GIT model achieves the best average performance across the three evaluated settings. We also observe that specialization consistently improves performance across different graphs, aligning with our theoretical analysis.
Comparison to Domain Experts. In addition to general GNN baselines applicable across various graphs, we compare our GIT model to domain experts specifically designed for molecules, including KVPLM (Zeng et al., 2022), MoMu (Su et al., 2022), Galactica (Taylor et al., 2022), and the recent SOTA model, GIMLET (Zhao et al., 2023). The results are presented in Table 23. We find that the general model pretrained on large-scale graphs generally underperforms compared to these domain experts. However, after specialization, the specialized model surpasses 3 out of 4 domain experts on average and outperforms the best expert model, GIMLET, on 4 out of 8 datasets. This observation demonstrates that post-training the general model with a reasonable number of domain-specific instances can enable it to match or even surpass expert models designed for that domain. These results strongly support the effectiveness of task-trees in designing graph foundation models.
F.6 Domain: Temporal E-Commerce
@@
F.7 General Reasoning Capability of Specialized Models

### v2

@@ 5.5 Comparison Between Task-Trees and Subgraphs
We evaluate the efficiency and performance of subgraphs and task-trees in the fine-tuning setting, as shown in Figure 4 and Table 4. To ensure a fair comparison, we implement a subgraph-based variant of GIT (GIT-SubG) by replacing task-trees with subgraphs. The results demonstrate that task-trees achieve both higher efficiency and improved model performance. While Table 4 reports results from three domains, the evaluation spans multiple tasks, including node classification (academic networks), edge classification (knowledge graphs), and graph classification (molecular networks).
6 Conclusion and Limitations
We introduce task-trees as basic learning instances for aligning graph-based tasks and provide both theoretical and empirical analyses to demonstrate their effectiveness. Building on task-trees, we develop GIT, a pretrained model that leverages a small set of graphs to enhance performance across more than 30 graphs spanning five domains. While our findings validate that task-trees preserve cross-task generalities at least partially, identifying the exact transferable patterns shared across graphs remains an open challenge. Further investigation from both empirical and theoretical perspectives is a promising direction for future research.
Impact Statement
@@ 6 Conclusion and Limitations
From an industry perspective, we offer GIT as a foundational tool for graph-structured data. Additionally, since GIT can be quickly adapted to specific domains, we hope it will support applications where label acquisition is difficult and model training is time-consuming. There is no clear ethical considerations on the model architecture itself.
@@ Appendix A Related Work
Graph Foundation Models. Foundation models are designed as general-purpose solvers capable of handling various tasks across different domains. For instance, LLMs, the foundation models in natural language processing, are capable of performing tasks such as summarization, translation, and entity recognition, as well as question-answering. However, building such versatile foundation models for graphs presents unique challenges due to the inherent feature, structural, and task heterogeneity across different graph domains and tasks. To address these challenges, Qiu et al. (2020) pretrained GNNs using subgraphs as basic units, mitigating structural heterogeneity. Building on this, Sun et al. (2023) reformulated node-, edge-, and graph-level tasks into subgraph-level tasks, tackling task heterogeneity. Additionally, Huang et al. (2023) and Liu et al. (2024) applied LLMs to unify the feature spaces of cross-domain graphs, addressing feature heterogeneity. These approaches enable models to operate on cross-domain and cross-task graphs. Further advancements, such as He & Hooi (2024) and Li et al. (2024), improve node embeddings by jointly optimizing GNN and LLM encoders, facilitating various downstream tasks like few-shot learning and zero-shot learning. However, most of these approaches rely on subgraphs as the primary learning instances, which can result in inefficient training and reduced expressiveness, as discussed in the main paper. Other efforts to resolve feature heterogeneity include methods like singular vector decomposition (SVD) (Zhao et al., 2024a; Yu et al., 2024), non-parametric encoders (Zhao et al., 2024b), or synthesizing text-attributed graphs (Wang et al., 2025). Notably, a recent study by Wang et al. (2024b) explores computation trees as transferable patterns in graphs. However, our work differs in three key aspects: (1) While Wang et al. (2024b) focus on identifying transferable patterns across graphs, our approach is centered on designing fundamental learning instances. (2) Our theoretical framework also serves as a foundational basis for Wang et al. (2024b). (3) Our proposed GIT significantly simplifies the model proposed by Wang et al. (2024b), demonstrating a minimally applicable design that retains effectiveness.
@@
Appendix B Additional Model Discussion
B.1 Why Does the General Model Need Specialization?
@@
B.2 More Analysis on Domain Regularizer
@@
B.3 Discussion on Homophily and Heterophily
@@
B.4 Discussion on Model Expressiveness
@@
B.5 Scaling Law
Model Size. We evaluated the model’s performance with different hidden dimensions, with results by domain presented in Figure 7. The results cover both basic fine-tuning and in-context learning, and comprehensive details are provided in Appendix K. We observe that increasing the number of hidden dimensions from 128 to 2,048 significantly improves model performance across all domains. We hypothesize that this improvement is due to the additional parameters, which enhance the model’s ability to memorize shared patterns across graphs. The observation indicates the potential existence of scaling laws when using task-trees as the basic learning instances.
@@ B.5 Scaling Law
B.6 Limitations and Future Work.
We treat task-trees as basic learning instances in graphs, similar to how images are used in vision and text in language. Despite we have validated that cross-graph generalities are, at least partially, preserved in task-trees, identifying the exact transferable patterns shared across graphs remains a challenge. Further exploration of this issue from both empirical and theoretical perspectives is worthwhile. Additionally, for message-passing GNNs, converting graphs into tree structures inevitably leads to some loss of information, limiting the model expressiveness. Advanced GNNs (Morris et al., 2019, 2023) or techniques (Zhang et al., 2021) could be used to enhance expressiveness. Moreover, as GIT is a relatively simple model, investigating potential extensions is also a promising direction, as discussed in Appendix C.
Appendix C Potential Model Extensions
C.1 Pretraining
C.2 Specialization via Instruction Tuning
How to Define Instructions? In this paper, as we focus on experiments with text-attributed graphs, we define instructions as label descriptions encoded by LLMs. However, this approach is not applicable to non-textual graphs. Other methods could be explored to define instructions, such as using proxy models (Hu et al., 2019) or graph heuristics (Jin et al., 2020) to generate instructions.
How to Leverage SFT Data? In scenarios with limited instructions, standard supervised fine-tuning may struggle to capture sufficient knowledge of the target domain. To address this, methods could be proposed to better utilize the unlabeled instances in the SFT dataset, thus enhancing model adaptation (Sohn et al., 2020).
C.3 More Scenarios.
The paper leverages text-attributed graphs to align node features. However, the pre-processing of TAGs can be time-consuming, raising the challenge of how to effectively apply the model to graphs without aligned node features. Furthermore, while we primarily focus on homogeneous graphs in this work, most real-world applications involve heterogeneous graphs. Addressing the question of how to design a single model capable of handling various types of graphs remains an open challenge. Finally, applying the model to specific applications Zhang et al. (2024c), which may exhibit unique characteristics, is another important consideration for future research.
@@ Appendix D Proof
By using a formulation like expression 11, we can decompose the impact of different layers, facilitating further analysis of the upper bound on the distance between two task-trees. Next, we will analyze the upper bound of each term. To begin, we first introduce a lemma.
@@ E.5 Hyper-Parameters
GIT. The model architecture and pretraining parameters of our GIT are presented in Table 7. The specific fine-tuning hyperparameters, categorized by domain, are shown in Tables 8, 9, 10, 11, and 12. For in-context learning and zero-shot learning results without fine-tuning, the general model does not involve any hyperparameters. For the specialized model, we tune the hyperparameters of SFT epochs from 10 to 500, in steps of 10, the SFT learning rate from 1e-4, 1e-5, 1e-6, 1e-7, 1e-8, and the normalization method from None, BN.
Appendix F Comprehensive Model Results
F.1 Domain: Academia
@@
F.2 Domain: E-Commerce
@@
F.3 Domain: Knowledge Base
@@
F.4 Domain: Molecule
Graph Classification. We evaluate fine-tuning, in-context learning, and zero-shot learning in this domain. The fine-tuning and in-context learning settings are consistent with those used in previous domains. For zero-shot learning, however, we follow the approach of Zhao et al. (2023) by assessing zero-shot performance on the original test set. The graph classification results are presented in Table 18. Our GIT model achieves the best average performance across the three evaluated settings. We also observe that specialization consistently improves performance across different graphs, aligning with our theoretical analysis.
@@ F.4 Domain: Molecule
F.5 Domain: Temporal E-Commerce
@@
Appendix G Comparison to Domain Experts
G.1 Domain: Knowledge Base
In addition to comparing GIT to standard baselines applicable for all graphs, we also evaluate it against Ultra, a foundation model specifically designed for knowledge graphs. As a domain expert, Ultra is compared to Expert GIT (pretrained on all KGs) and Specialized GIT (pretrained on default datasets and fine-tuned on FB15K237), with the results presented in Table 21. We find that the two domain experts, Ultra and Expert GIT, achieve comparable performance, though Ultra significantly outperforms Expert GIT in certain settings. This may be due to Ultra learning more fine-grained relational information within KGs. Notably, Specialized GIT also performs comparably to both domain experts, highlighting the potential of specialization. We believe this is because the distributions of KGs are more similar to each other compared to graphs from other domains.
3-shot Finetune
Ultra Expert GIT Specialized GIT Ultra Expert GIT Specialized GIT
Average 68.53 67.72 67.80 76.26 76.03 76.15
G.2 Domain: Molecule
In addition to general GNN baselines applicable across various graphs, we compare our GIT model to domain experts specifically designed for molecules, including KVPLM (Zeng et al., 2022), MoMu (Su et al., 2022), Galactica (Taylor et al., 2022), and the recent SOTA model, GIMLET (Zhao et al., 2023). The results are presented in Table 22. We find that the general model pretrained on large-scale graphs generally underperforms compared to these domain experts. However, after specialization, the specialized model surpasses 3 out of 4 domain experts on average and outperforms the best expert model, GIMLET, on 4 out of 8 datasets. This observation demonstrates that post-training the general model with a reasonable number of domain-specific instances can enable it to match or even surpass expert models designed for that domain. These results strongly support the effectiveness of task-trees in designing graph foundation models.
HIV BBBP BACE TOXCAST CYP450 TOX21 MUV PCBA Avg.
KVPLM 61.20 60.20 51.26 50.96 59.22 49.17 61.72 48.11 55.23
MoMu 50.26 49.81 66.56 52.38 57.98 57.57 60.51 51.50 55.82
Galactica-1.3B 33.85 53.94 56.48 51.23 46.86 49.46 57.15 52.02 50.12
GIMLET 66.24 59.39 69.57 59.04 71.25 61.19 64.39 62.11 64.15
GIT - G 56.76 54.76 33.66 51.55 63.21 56.83 53.71 56.25 53.34
GIT - S 66.14 62.16 52.27 58.30 69.75 63.45 65.32 65.26 62.83
indicates the results from paper (Zhao et al., 2023).
Appendix H Ablation on Specialization (SFT)
Table 23: The impact of SFT datasets in zero-shot setting.
SFT Data HIV BBBP BACE TOXCAST CYP450 TOX21 MUV PCBA Avg.
PCBA 66.14 62.16 52.27 58.30 69.75 63.45 65.32 65.26 62.83
HIV 66.28 45.97 43.35 52.78 64.50 57.86 53.46 46.57 53.85
Appendix I Efficiency Analysis on Industry-scale Data
In order to further evaluate the efficiency of task-trees over subgraphs, we conduct experiments on a synthetic graph, termed as ogbn-products x50. This dataset is derived from the existing ogbn-products graph, which contains 316,513 nodes and 19,337,745 edges (we use a subset of original ogbn-products in our experiments). Specifically, we replicate the graph 50 times, reindex the nodes in each replicated graph, and combine them into a single dataset. This synthetic dataset contains 15,825,650 nodes (there are 111,059,956 nodes in ogbn-papers100M) and 966,887,250 edges (there are 1,615,685,872 edges in ogbn-papers100M). To ensure connectivity among these replicated graphs, we randomly flip 5% of the edges to connect the individual components, creating a more realistic large-scale graph structure.
We evaluated the time consumption per epoch on the ogbn-products x50 dataset and compared the efficiency of our proposed method, GIT-task-tree, against GIT-SubG (a variation that replaces task-trees with subgraphs). With a batch size of 512—selected to avoid out-of-memory issues in subgraph-based methods—the results are as follows:
Table 24: Efficiency comparison between subgraphs and task-trees in ogbn-products x50 (15,825,650 nodes and 966,887,250 edges).
GIT-SubG GIT-Task-Tree
Sampling 546s 546s
Subgraph/Tree Extraction 948s 0s
Encoding 5,717s 2,429s
All 7,211s 2,975s
These results underscore the significant efficiency advantages of our method. Notably, the efficiency gap between GIT-task-tree and GIT-SubG on this dataset is larger than what we observed during pretraining (i.e., 208 seconds per epoch for GIT-task-tree versus 280 seconds per epoch for GIT-SubG). We attribute this to the substantial overlap between subgraphs in this dataset, which increases the space and time required for processing.
Appendix J General Reasoning Capability of Specialized Models